OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5).

OpenStudy (anonymous):

A.-1 B.1 C.0 D.1/2

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

Slope =-1?

pooja195 (pooja195):

\[\LARGE \frac{-5-(-3)}{3-1}=slope\] \[\LARGE \frac{-5+3}{3-1}=slope\]

pooja195 (pooja195):

-1 is correct good job :)

OpenStudy (anonymous):

THANK YOU

pooja195 (pooja195):

:)